OpenStudy (anonymous):

how do youu solve for x in 2/3x + 1/2=5/6 ?

OpenStudy (gw2011):

Multiply each side by 6 in order to eliminate the denominator and you get: 4x + 3 = 5 4x = 5 - 3 4x = 2 x = 2/4 = 1/2